### v1

@@ 1 Introduction
Recent advances in generative modeling—particularly diffusion models [1, 2], and autoregressive methods [3, 4]—have significantly improved the quality of image [5, 6], video [7, 8], and 3D content synthesis [9, 10]. Despite this progress, most generative models are trained using surrogate objectives such as maximum likelihood estimation or reconstruction loss, which are often misaligned with human perception [11, 12], semantics [13], or physical plausibility [14]. As a result, generated outputs can exhibit artifacts such as motion inconsistency [15, 16], or structural errors [17].
@@
This survey presents a comprehensive overview of how reinforcement learning is being integrated into modern generative modeling across multiple domains. We begin with the conceptual evolution of RL—from its origins in solving Markov decision processes to its emergence as a general-purpose optimization framework embedded within broader learning systems. We then examine its application in image generation, where RL techniques are used to improve semantic alignment, aesthetic quality, and controllability through policy optimization and preference-based learning. In the video domain, RL contributes to efficient sampling, temporal coherence, physical plausibility, and alignment with user intent. Finally, we explore the use of RL in 3D content generation, including point cloud reconstruction, mesh synthesis, multi-view consistency, and human motion modeling. In all, these threads illustrate the growing role of RL as a versatile and principled mechanism for guiding generative models toward complex, structured, and human-aligned objectives.
Across these domains, RL serves as a flexible and interpretable optimization layer that complements existing architectures. By reframing generation as an interactive process driven by evaluative feedback, RL allows generative models to move beyond static supervision and toward learning from preferences, outcomes, or real-world constraints. This survey aims to map the methodological landscape, identify cross-domain patterns, and offer insights into future research directions at the intersection of reinforcement learning and generative modeling.
@@ 2 RL revolution
2.1 Reinforcement Learning as MDP Solvers
The first phase of reinforcement learning was grounded in the formalism of Markov Decision Processes (MDPs), which define sequential decision-making under uncertainty through a set of states, actions, transition dynamics, and reward functions [37, 38, 49]. The goal was to compute an optimal policy that maximizes expected cumulative reward. Early RL efforts closely mirrored the dynamic programming solutions for MDPs, assuming either full or sampled access to the environment.
@@ 2.2 Emergence of Specialized Subfields
Offline reinforcement learning, model-based reinforcement learning, and other subfields (multi-agent RL, safe RL, etc.) emerged as semi-independent research areas. Each introduced new assumptions or additional components to the RL paradigm to tackle specific challenges: removing the need for active environment interaction, incorporating knowledge of dynamics, handling multiple agents or risk, and so on. This fragmentation represented a paradigm shift: rather than viewing RL as just an agent solving an MDP, RL could be re-formulated to fit constraints of data availability, environment knowledge, or task structure.
Offline Reinforcement Learning: Offline reinforcement learning addresses the challenge of learning effective policies from fixed datasets without any further environment interaction. This setting is particularly valuable in domains where exploration is costly, risky, or infeasible—such as autonomous driving, healthcare, and industrial control. By reframing RL as a static-data learning problem, offline RL brings it closer to supervised learning, enabling more scalable and safer deployment in real-world systems.
@@
Multi-Agent Reinforcement Learning: When multiple agents interact in the same environment, the single-agent MDP framework no longer holds because the environment becomes non-stationary from each agent’s perspective. This challenge led to the development of multi-agent reinforcement learning (MARL) [68, 69, 70, 71], which addresses issues such as non-stationarity, cooperation, competition, and equilibrium strategies. Key contributions include MADDPG [72], which introduced centralized training with decentralized execution, allowing agents to condition their critics on global information during training while executing independently at test time. Other approaches, such as LOLA [73], incorporated opponent modeling by anticipating how an agent’s updates influence others’ learning. MARL expanded the classical RL framework to stochastic games and multi-agent MDPs, drawing on concepts from game theory to develop stable learning dynamics and policy solutions.
Risk-Sensitive and Robust Reinforcement Learning: Traditional RL optimizes the expected return, which may overlook performance variability or safety-critical failures. Risk-sensitive methods [74] address this by optimizing alternative criteria, such as minimizing variance or ensuring favorable worst-case outcomes. Distributional RL models [75, 76] the entire return distribution rather than its mean, enabling richer decision-making. Robust RL [77, 78, 79] further addresses uncertainty in dynamics or rewards, seeking policies that perform well under perturbations or adversarial changes. Techniques such as domain randomization [80] and adversarial training [79] have been used to enhance robustness and stress-test policies. These approaches reflect a shift from reward maximization toward more reliable and resilient objectives.
Safe Reinforcement Learning: Safe RL focuses on satisfying explicit safety constraints throughout training and deployment. It often formulates the problem as a constrained MDP, incorporating methods such as Lagrangian optimization [81], reward shaping [82], or safety shields [83, 84] to enforce limits on unsafe behavior. This is particularly critical in domains like autonomous driving [85] and robotics [86], where unsafe exploration can lead to catastrophic failures. Safe RL also intersects with exploration research, aiming to develop learning strategies that remain within safety boundaries while still enabling policy improvement.
@@ 2.3 Redefining the Problem — RL as Learning to Simulate, Optimize, and Align
During the late 2010s to mid-2020s, reinforcement learning evolved from optimizing fixed reward functions in known environments to serving as a modular component within broader learning systems. This phase emphasized two key directions: first, using RL to optimize over models learned from data, such as simulators or reward functions [44, 43]; second, aligning agent behavior with nuanced human goals that cannot be easily captured by scalar rewards [33, 45]. The focus shifted from refining policy optimization algorithms to designing the outer structure of the learning task itself, including how preferences are modeled and how environment dynamics are inferred [87, 88]. RL became less about the optimizer and more about how the problem is formulated.
Reinforcement Learning from Human Feedback: Reinforcement Learning from Human Feedback (RLHF) [33, 45] has become a high-impact application of reinforcement learning, particularly in aligning large language models with human intentions. Rather than relying on hand-crafted reward functions, RLHF uses human preferences to define the objective. The typical pipeline involves collecting human demonstrations or rankings of model outputs [33, 87], training a reward model to predict these preferences, and then fine-tuning the model using RL to maximize this learned reward [45]. This transforms vague goals like “be helpful” into a concrete RL problem: the model generates responses, the reward is given by the preference model, and RL optimizes accordingly.
The success of InstructGPT [45] demonstrated the effectiveness of this approach. A 1.3B parameter model fine-tuned via RLHF outperformed the original 175B GPT-3 in instruction following, while reducing harmful or irrelevant outputs. RLHF builds on earlier work such as preference-based RL [33], which showed that reward functions can be learned from human comparisons rather than manually specified. In high-dimensional spaces like language or vision, such feedback provides a scalable supervision signal. Importantly, the core challenge in RLHF lies not in the RL algorithm itself, but in designing robust reward models from subjective data [88]. This paradigm has since extended to domains like image generation and robotics, where human-aligned behavior is critical [87]. Overall, RLHF exemplifies a broader view of reinforcement learning—as a flexible framework for aligning complex models with implicit goals through interaction and evaluative feedback.
World Models and the Resurgence of Model-Based Reinforcement Learning: Model-based reinforcement learning has re-emerged as a promising solution to the sample inefficiency of traditional RL [48, 42]. By learning a model of environment dynamics from data, agents can simulate interactions internally, enabling policy training with significantly fewer real-world interactions. Early work such as World Models and PETS demonstrated that policies trained in learned simulators could transfer successfully to real environments [48, 42], highlighting the potential of model-based approaches.
@@
Recent architectures such as DreamerV3 and Diffusion-QL integrate world models with generative and preference-aware objectives, illustrating a shift in focus from algorithmic innovation to task specification and system-level design [90, 91]. In this view, reinforcement learning increasingly involves first learning the underlying Markov decision process—its dynamics and rewards—followed by planning or policy optimization within that learned model. This two-stage formulation offers greater flexibility and data efficiency, and has proven effective even when the learned model is imperfect. The result is a broader, more adaptable vision of RL, where modeling, simulation, and optimization are deeply intertwined.
2.4 Reinforcement Learning as a General Substrate for Decision-Making
Recently, reinforcement learning is increasingly viewed not as a standalone technique but as a general-purpose framework for decision-making across diverse domains. Rather than focusing solely on interactive agents in fixed environments, modern RL is embedded into larger systems involving planning, simulation, supervised learning, and human feedback.
A key feature of this phase is the emphasis on problem formulation over algorithmic novelty. In applications like RLHF [45], the choice of RL algorithm (e.g., PPO [92]) is secondary; the critical component is the learned reward model derived from human preferences. Similarly, in systems like AlphaGo [93], RL serves as one module within a broader architecture that includes imitation, self-play, and tree search. In these cases, RL acts as a fine-tuning layer that optimizes a well-defined objective.
@@ 3 RL for Image Generation
Table 1: Comprehensive overview of reinforcement learning methods applied to image generation, organized chronologically.
@@ 3.1 PPO-based Image Generation
PPO-based image generation represents an early exploration of reinforcement learning for improving image synthesis, particularly in aligning model outputs with human preferences, enhancing semantic consistency, and enabling controllable generation. These approaches treat image generation, especially in diffusion or autoregressive model, as a sequential decision-making process and leverage learned reward signals to guide policy optimization.
One representative approach is Denoising Diffusion Policy Optimization (DDPO) [96], which formulates the denoising process in diffusion models as a multi-step MDP. This allows the model to directly optimize downstream generation objectives, rather than relying solely on likelihood-based training. Concurrently, DPOK [94] also advances this direction by framing the fine-tuning of text-to-image diffusion models as an online Reinforcement Learning problem within a multi-step Markov Decision Process (MDP) framework. It optimizes feedback-trained rewards using policy gradient, critically incorporating Kullback–Leibler (KL) regularization to concurrently enhance both image quality and text-image alignment.
@@
Beyond generation-stage optimization, PPO-based methods have also been applied to prompt engineering. Promptist [95] proposes a two-stage pipeline that combines supervised fine-tuning (SFT) with PPO-based reinforcement learning to automatically optimize user prompts. This improves alignment between user intent and the model’s internal preference signals. POSI [101] extends this idea to safe text-to-image generation by using a custom reward function that balances semantic alignment with safety constraints. PAE [103] further introduces dynamic prompt optimization, enabling fine-grained control over prompt weights and injection timing. By integrating SFT with online PPO, PAE optimizes for aesthetics, semantic fidelity, and user preference under a unified reward formulation.
@@ 3.3 GRPO-based Image Generation
Generalized Reinforcement Policy Optimization (GRPO) is a recent extension of policy optimization that offers improved generalization and training stability. As a variant of PPO, GRPO has gained traction in visual content generation and has been applied across diffusion- and flow-based pipelines.
@@
Flow-GRPO [118] extends GRPO to flow-based generative models by reformulating generation as a stochastic differential equation (SDE). This allows for effective reinforcement learning-based exploration. It also introduces a denoising-step shrinking strategy to improve both sample efficiency and output fidelity.
@@ 5 RL for 3D Content Generation
Figure 2: This figure presents the proportional breakdown of recent research topics applying reinforcement learning to 3D generation, including Text-to-NeRF/3D Gaussian Splatting, 3D diffusion models, multi-view consistency optimization, human motion synthesis, point cloud modeling, and others. The balanced distribution across these areas reflects a rapidly emerging and diversifying field, where no single paradigm dominates. This suggests that RL is being broadly explored as a general-purpose optimization tool across the full spectrum of 3D generation tasks.
Text- and Image-Conditioned 3D Generation: Recent advances in 3D generation have increasingly explored the use of reinforcement learning to enhance controllability, geometric consistency, and human preference alignment. In particular, RL provides a natural framework for optimizing non-differentiable objectives, enabling fine-grained reward-driven adaptation of 3D assets generated from text or image prompts.
@@ 5.2 Multi-View Model Enhancement
Carve3D [156] introduces Reinforcement Learning Fine-Tuning (RLFT) into multi-view diffusion models. Starting from a pretrained text-to-image model, Carve3D applies RLFT using a novel Multi-view Reconstruction Consistency (MRC) metric. This metric compares images generated from multiple viewpoints with views rendered from a NeRF reconstructed by a large sparse-view model (LRM). The MRC score serves as the reward signal, guiding the diffusion model to produce outputs that are both visually plausible and geometrically consistent across views.
@@ 5.3 Other Domain-Specific 3D Applications
Reinforcement learning has not only been adopted to improve the quality and alignment of general 3D assets, but also extended to domain-specific generation tasks that require structured control, physical reasoning, or multi-object coordination. These applications further exemplify the versatility of RL as a general decision-making substrate for visual content generation.
@@
Table 4: Comprehensive overview of reinforcement learning methods applied to 3D generation.
@@ 6.3 Future Work
The convergence of reinforcement learning and generative modeling is reshaping how we think about generation—not as a static mapping from input to output, but as an interactive, iterative, and goal-driven process. As generative systems become more autonomous and user-facing, the ability to learn from feedback, optimize complex objectives, and adapt to diverse preferences will be indispensable. Reinforcement learning provides the theoretical tools and algorithmic foundations to meet these demands.

### v2

@@ 1 Introduction
Recent advances in generative modeling—particularly diffusion models [1, 2] and autoregressive approaches [3, 4]—have substantially enhanced the quality of synthesized content across image [5, 6], video [7, 8], and 3D contents [9, 10]. These breakthroughs have increasingly demonstrated the potential to extend beyond content generation into broader AI paradigms [11, 12, 13, 14, 15, 16, 17, 18, 19]. Despite this progress, most generative models are trained using surrogate objectives such as maximum likelihood estimation or reconstruction loss, which are often misaligned with human perception [20, 21], semantics [22], or physical plausibility [23]. As a result, generated outputs can exhibit artifacts such as motion inconsistency [24, 25], or structural errors [26].
@@
This survey presents a comprehensive overview of how RL is being integrated into modern generative modeling across multiple domains. We begin with the conceptual evolution of RL—from its origins in solving Markov decision processes to its emergence as a general-purpose optimization framework embedded within broader learning systems. We then examine its application in image generation, where RL techniques are used to improve semantic alignment, aesthetic quality, and controllability through policy optimization and preference-based learning. In the video domain, RL contributes to efficient sampling, temporal coherence, physical plausibility, and alignment with user intent. Finally, we explore the use of RL in 3D content generation, including point cloud reconstruction, mesh synthesis, multi-view consistency, and human motion modeling. In all, these threads illustrate the growing role of RL as a versatile and principled mechanism for guiding generative models toward complex, structured, and human-aligned objectives.
Across these domains, RL serves as a flexible and interpretable optimization layer that complements existing architectures. By reframing generation as an interactive process driven by evaluative feedback, RL allows generative models to move beyond static supervision and toward learning from preferences, outcomes, or real-world constraints. This survey aims to map the methodological landscape, identify cross-domain patterns, and offer insights into future research directions at the intersection of RL and generative modeling.
@@ 2 RL Revolution
2.1 RL as MDP Solvers
The first phase of RL was grounded in the formalism of Markov Decision Processes (MDPs), which define sequential decision-making under uncertainty through a set of states, actions, transition dynamics, and reward functions [46, 47, 58]. The goal was to compute an optimal policy that maximizes expected cumulative reward. Early RL efforts closely mirrored the dynamic programming solutions for MDPs, assuming either full or sampled access to the environment.
@@ 2.2 Emergence of Specialized Subfields
Offline RL, model-based RL, and other subfields (multi-agent RL, safe RL, etc.) emerged as semi-independent research areas. Each introduced new assumptions or additional components to the RL paradigm to tackle specific challenges: removing the need for active environment interaction, incorporating knowledge of dynamics, handling multiple agents or risk, and so on. This fragmentation represented a paradigm shift: rather than viewing RL as just an agent solving an MDP, RL could be re-formulated to fit constraints of data availability, environment knowledge, or task structure.
Offline RL: Offline RL addresses the challenge of learning effective policies from fixed datasets without any further environment interaction. This setting is particularly valuable in domains where exploration is costly, risky, or infeasible—such as autonomous driving, healthcare, and industrial control. By reframing RL as a static-data learning problem, offline RL brings it closer to supervised learning, enabling more scalable and safer deployment in real-world systems.
@@
Multi-Agent RL: When multiple agents interact in the same environment, the single-agent MDP framework no longer holds because the environment becomes non-stationary from each agent’s perspective. This challenge led to the development of multi-agent RL (MARL) [75, 76, 77, 78], which addresses issues such as non-stationarity, cooperation, competition, and equilibrium strategies. Key contributions include MADDPG [79], which introduced centralized training with decentralized execution, allowing agents to condition their critics on global information during training while executing independently at test time. Other approaches, such as LOLA [80], incorporated opponent modeling by anticipating how an agent’s updates influence others’ learning. MARL expanded the classical RL framework to stochastic games and multi-agent MDPs, drawing on concepts from game theory to develop stable learning dynamics and policy solutions.
Risk-Sensitive and Robust RL: Traditional RL optimizes the expected return, which may overlook performance variability or safety-critical failures. Risk-sensitive methods [81] address this by optimizing alternative criteria, such as minimizing variance or ensuring favorable worst-case outcomes. Distributional RL models [82, 83] the entire return distribution rather than its mean, enabling richer decision-making. Robust RL [84, 85, 86] further addresses uncertainty in dynamics or rewards, seeking policies that perform well under perturbations or adversarial changes. Techniques such as domain randomization [87] and adversarial training [86] have been used to enhance robustness and stress-test policies. These approaches reflect a shift from reward maximization toward more reliable and resilient objectives.
Safe RL: Safe RL focuses on satisfying explicit safety constraints throughout training and deployment. It often formulates the problem as a constrained MDP, incorporating methods such as Lagrangian optimization [88], reward shaping [89], or safety shields [90, 91] to enforce limits on unsafe behavior. This is particularly critical in domains like autonomous driving [92] and robotics [93], where unsafe exploration can lead to catastrophic failures. Safe RL also intersects with exploration research, aiming to develop learning strategies that remain within safety boundaries while still enabling policy improvement.
@@ 2.3 Redefining the Problem — RL as Learning to Simulate, Optimize, and Align
During the late 2010s to mid-2020s, RL evolved from optimizing fixed reward functions in known environments to serving as a modular component within broader learning systems. This phase emphasized two key directions: first, using RL to optimize over models learned from data, such as simulators or reward functions [53, 52]; second, aligning agent behavior with nuanced human goals that cannot be easily captured by scalar rewards [42, 54]. The focus shifted from refining policy optimization algorithms to designing the outer structure of the learning task itself, including how preferences are modeled and how environment dynamics are inferred [94, 95]. RL became less about the optimizer and more about how the problem is formulated.
RL from Human Feedback: RL from Human Feedback (RLHF) [42, 54] has become a high-impact application of RL, particularly in aligning large language models with human intentions. Rather than relying on hand-crafted reward functions, RLHF uses human preferences to define the objective. The typical pipeline involves collecting human demonstrations or rankings of model outputs [42, 94], training a reward model to predict these preferences, and then fine-tuning the model using RL to maximize this learned reward [54]. This transforms vague goals like “be helpful” into a concrete RL problem: the model generates responses, the reward is given by the preference model, and RL optimizes accordingly.
The success of InstructGPT [54] demonstrated the effectiveness of this approach. A 1.3B parameter model fine-tuned via RLHF outperformed the original 175B GPT-3 in instruction following, while reducing harmful or irrelevant outputs. RLHF builds on earlier work such as preference-based RL [42], which showed that reward functions can be learned from human comparisons rather than manually specified. In high-dimensional spaces like language or vision, such feedback provides a scalable supervision signal. Importantly, the core challenge in RLHF lies not in the RL algorithm itself, but in designing robust reward models from subjective data [95]. This paradigm has since extended to domains like image generation and robotics, where human-aligned behavior is critical [94]. Overall, RLHF exemplifies a broader view of RL—as a flexible framework for aligning complex models with implicit goals through interaction and evaluative feedback.
World Models and the Resurgence of Model-Based RL: Model-based RL has re-emerged as a promising solution to the sample inefficiency of traditional RL [57, 51]. By learning a model of environment dynamics from data, agents can simulate interactions internally, enabling policy training with significantly fewer real-world interactions. Early work such as World Models and PETS demonstrated that policies trained in learned simulators could transfer successfully to real environments [57, 51], highlighting the potential of model-based approaches.
@@
Recent architectures such as DreamerV3 and Diffusion-QL integrate world models with generative and preference-aware objectives, illustrating a shift in focus from algorithmic innovation to task specification and system-level design [97, 98]. In this view, RL increasingly involves first learning the underlying Markov decision process—its dynamics and rewards—followed by planning or policy optimization within that learned model. This two-stage formulation offers greater flexibility and data efficiency, and has proven effective even when the learned model is imperfect. The result is a broader, more adaptable vision of RL, where modeling, simulation, and optimization are deeply intertwined.
2.4 RL as a General Substrate for Decision-Making
Recently, RL is increasingly viewed not as a standalone technique but as a general-purpose framework for decision-making across diverse domains. Rather than focusing solely on interactive agents in fixed environments, modern RL is embedded into larger systems involving planning, simulation, supervised learning, and human feedback.
A key feature of this phase is the emphasis on problem formulation over algorithmic novelty. In applications like RLHF [54], the choice of RL algorithm (e.g., PPO [99], GRPO [100, 101]) is secondary; the critical component is the learned reward model derived from human preferences. Similarly, in systems like AlphaGo [102], RL serves as one module within a broader architecture that includes imitation, self-play, and tree search. In these cases, RL acts as a fine-tuning layer that optimizes a well-defined objective.
@@ 3 RL for Image Generation
Table 1: Comprehensive overview of RL methods applied to image generation, organized chronologically.
@@ 3.1 PPO-based Image Generation
PPO-based image generation represents an early exploration of RL for improving image synthesis, particularly in aligning model outputs with human preferences, enhancing semantic consistency, and enabling controllable generation. These approaches treat image generation, especially in diffusion or autoregressive model, as a sequential decision-making process and leverage learned reward signals to guide policy optimization.
One representative approach is Denoising Diffusion Policy Optimization (DDPO) [105], which formulates the denoising process in diffusion models as a multi-step MDP. This allows the model to directly optimize downstream generation objectives, rather than relying solely on likelihood-based training. Concurrently, DPOK [103] also advances this direction by framing the fine-tuning of text-to-image diffusion models as an online RL problem within a multi-step Markov Decision Process (MDP) framework. It optimizes feedback-trained rewards using policy gradient, critically incorporating Kullback–Leibler (KL) regularization to concurrently enhance both image quality and text-image alignment.
@@
Beyond generation-stage optimization, PPO-based methods have also been applied to prompt engineering. Promptist [104] proposes a two-stage pipeline that combines supervised fine-tuning (SFT) with PPO-based RL to automatically optimize user prompts. This improves alignment between user intent and the model’s internal preference signals. POSI [110] extends this idea to safe text-to-image generation by using a custom reward function that balances semantic alignment with safety constraints. PAE [112] further introduces dynamic prompt optimization, enabling fine-grained control over prompt weights and injection timing. By integrating SFT with online PPO, PAE optimizes for aesthetics, semantic fidelity, and user preference under a unified reward formulation.
@@ 3.3 GRPO-based Image Generation
Group Relative Policy Optimization (GRPO) [100] is a recent extension of policy optimization that offers improved generalization and training stability. As a variant of PPO, GRPO has gained traction in visual content generation and has been applied across diffusion- and flow-based pipelines.
@@
Flow-GRPO [126] extends GRPO to flow-based generative models by reformulating generation as a stochastic differential equation (SDE). This allows for effective RL-based exploration. It also introduces a denoising-step shrinking strategy to improve both sample efficiency and output fidelity.
@@ 5 RL for 3D Content Generation
Figure 2: This figure presents the proportional breakdown of recent research topics applying RL to 3D generation, including Text-to-NeRF/3D Gaussian Splatting, 3D diffusion models, multi-view consistency optimization, human motion synthesis, point cloud modeling, and others. The balanced distribution across these areas reflects a rapidly emerging and diversifying field, where no single paradigm dominates. This suggests that RL is being broadly explored as a general-purpose optimization tool across the full spectrum of 3D generation tasks.
Text- and Image-Conditioned 3D Generation: Recent advances in 3D generation have increasingly explored the use of RL to enhance controllability, geometric consistency, and human preference alignment. In particular, RL provides a natural framework for optimizing non-differentiable objectives, enabling fine-grained reward-driven adaptation of 3D assets generated from text or image prompts.
@@ 5.2 Multi-View Model Enhancement
Carve3D [164] introduces RL Fine-Tuning (RLFT) into multi-view diffusion models. Starting from a pretrained text-to-image model, Carve3D applies RLFT using a novel Multi-view Reconstruction Consistency (MRC) metric. This metric compares images generated from multiple viewpoints with views rendered from a NeRF reconstructed by a large sparse-view model (LRM). The MRC score serves as the reward signal, guiding the diffusion model to produce outputs that are both visually plausible and geometrically consistent across views.
@@ 5.3 Other Domain-Specific 3D Applications
RL has not only been adopted to improve the quality and alignment of general 3D assets, but also extended to domain-specific generation tasks that require structured control, physical reasoning, or multi-object coordination. These applications further exemplify the versatility of RL as a general decision-making substrate for visual content generation.
@@
Table 4: Comprehensive overview of RL methods applied to 3D generation.
@@ 6.3 Future Work
The convergence of RL and generative modeling is reshaping how we think about generation—not as a static mapping from input to output, but as an interactive, iterative, and goal-driven process. As generative systems become more autonomous and user-facing, the ability to learn from feedback, optimize complex objectives, and adapt to diverse preferences will be indispensable. RL provides the theoretical tools and algorithmic foundations to meet these demands.